OpenStudy (anonymous):

Help me, determine whether each ordered pair is a solution of the given equation. Medal will be given. Y=2x^2;(1,2), (3,18)

OpenStudy (lucaz):

you can find using the values of x, the pair (1,2) 1=x and 2=y, plug in the x in the eqution: y=2(1)^2 you get y=2(1)=2 so (1,2) is a solution ofr this equation. do the same thing with the other pair